### v3

@@ 2.3 LLM Embeddings Achieve Strong Performance in Low-Data Regimes
| GTE-Qwen2-7B | |||||
| generic instructions | |||||
| no instructions | |||||
| no demographics | |||||
| no aggregated | |||||
| no visits (both) | |||||
| no conditions | |||||
| no medications | |||||
| no procedures | |||||
| LLM2Vec-Llama-3.1-8B | |||||
| generic instructions | |||||
| no instructions | |||||
| no demographics | |||||
| no aggregated | |||||
| no visits (both) | |||||
| no conditions | |||||
| no medications | |||||
| no procedures |
2.4 Effect of Serialization Components on LLM-Based EHR Embedding Performance
To evaluate the contribution of individual serialization components, we conducted ablation studies for both LLM embedding models (see Table 5). Replacing task-specific instructions with a generic prompt led to a marked performance decline, reducing average AUROC from 0.755 (0.730-0.780) to 0.739 (0.713-0.765) for GTE-Qwen2-7B and from 0.734 (0.707-0.762) to 0.724 (0.696-0.752) for LLM2Vec-Llama-3.1-8B. Removing the instruction entirely caused a further drop for GTE-Qwen2-7B to 0.724 (0.698-0.749), whereas LLM2Vec-Llama-3.1-8B showed minimal change, suggesting greater reliance on instruction prompting by GTE-Qwen2-7B. The most significant drop was observed in lab result prediction tasks, highlighting the importance of guidance via task-specific instructions. Aggregated measurements (body metrics, vital signs, and lab values) were crucial, especially for lab test prediction, where their removal reduced AUROC from 0.865 (0.858-0.873) to 0.714 (0.701-0.726) for GTE-Qwen2-7B and from 0.778 (0.768-0.789) to 0.706 (0.693-0.719) for LLM2Vec-Llama-3.1-8B. This underscores the predictive value of recent lab results and their effective encoding by the models. For conditions like thrombocytopenia, hyperkalemia, and hyponatremia, the performance drop was especially pronounced without this information for GTE-Qwen2-7B (see Fig. 8). Interestingly, removing aggregated values slightly improved new diagnosis prediction for GTE-Qwen2-7B, suggesting a more focused input may benefit some tasks. Visit-level information had a strong influence on operational outcome tasks. Its removal led to AUROC reductions from 0.771 (0.747-0.795) to 0.752 (0.728-0.776) for GTE-Qwen2-7B and from 0.753 (0.727-0.778) to 0.734 (0.709-0.759) for LLM2Vec-Llama-3.1-8B. However, for LLM2Vec-Llama-3.1-8B, removing visit data unexpectedly improved lab result prediction, suggesting possible difficulty in filtering relevant signals from verbose input. Removing demographics, conditions, medications, and procedures had only minor effects, indicating these features were either redundant or contributed little predictive signal in isolation. While aggregated information and visit data were key drivers for lab and operational outcomes, prediction of new diagnoses and chest X-ray findings was less sensitive to individual serialization components. This suggests these tasks rely on broader contextual patterns rather than specific structured elements.
2.5 Effect of Temporal Scope and Context Length on Embedding Effectiveness
To assess model sensitivity to input length and recency, we performed ablations on temporal window, context size, and chunking using EHRSHOT data. Both LLM embedding models showed improved performance when restricted to more recent data, with the 1-month window outperforming longer histories. This suggests that recent clinical information may carry the most predictive signal, while older records may contribute less relevant or potentially distracting details (see Fig. 6 and Section A.2). The models differed in optimal context size: GTE-Qwen2-7B achieved the best performance at the maximum 4,096-token context, whereas LLM2Vec-Llama-3.1-8B performed best at 2,048 tokens (Fig. 6 and Table 10). Both models suffered large drops at 512 tokens, a common constraint in older language models [37]. To test whether models processed the entire input in context, we evaluated chunked context representations, where fixed-size segments (e.g., 512–2,048 tokens) of a 4,096-token input were embedded separately and averaged (see Table 11). GTE-Qwen2-7B showed little degradation from chunking, indicating that simple averaging can be sufficient to preserve performance even under token limits. In contrast, LLM2Vec-Llama-3.1-8B benefited from chunking, suggesting it struggles with longer contiguous sequences and prefers shorter inputs. These results highlight differences in model-specific sequence processing and suggest practical approaches for adapting to context length constraints. Full tables and additional details are available in Section A.3 and Section A.4.
@@ 3 Discussion
Our study shows that general-purpose LLM embedding models, pre-trained on large-scale natural language corpora, can be effectively repurposed for clinical prediction using EHR data. Models such as GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B outperformed a strong count-based baseline and, in several clinical domains, matched or surpassed the dedicated EHR foundation model CLMBR-T-Base [13, 31]. This is notable, as CLMBR-T-Base was trained on data from the same health system as EHRSHOT. The strong performance of the LLM-based encoders on the UKB dataset highlights their generalization capabilities. GTE-Qwen2-7B consistently outperformed the domain-specific foundation model with at least 16 training examples, while LLM2Vec-Llama-3.1-8B achieved similar results. These results are particularly notable given the challenges posed by EHR data, including limited size, heterogeneity, and domain shift. Across a broad set of tasks, including few-shot settings, our findings indicate that knowledge acquired during large-scale text pre-training enables LLMs to extract clinically meaningful patterns from serialized EHRs. Model performance improved with larger model sizes and more advanced architectures, suggesting further gains are likely with future LLM development. In contrast, domain-specific EHR models remain constrained by the scarcity and fragmentation of large, diverse clinical datasets. Overall, our findings support general-purpose LLMs as scalable, adaptable, and high-performing alternatives for EHR representation and clinical prediction.
Our experiments suggest that several key factors contribute to the success of LLM embedding models for encoding EHR data. First, providing clear, task-specific instructions in the serialized EHR text was crucial, guiding models to focus on clinically relevant sections and improving tasks requiring targeted attention. This takes advantage of the instruction-tuned nature of LLM embedding models [27, 30], enabling them to extract meaningful patterns from heterogeneous EHR inputs. Second, aggregating structured data, particularly recent lab results and visit-level information, substantially improved performance, especially for lab test and operational outcome prediction. In contrast, components like demographics, medications, and condition codes contributed less consistently to model accuracy. Context size was also critical: GTE-Qwen2-7B effectively handled up to 4,096 tokens, while both models performed poorly with very short or overly long inputs. Intermediate time windows, especially those covering one month prior to prediction, yielded optimal performance, suggesting that recent clinical data carries more predictive value than long-term history for many tasks. For instance, lab values reflect dynamic physiological states and may lose relevance outside of a narrow temporal window. This indicates that LLM embeddings are particularly effective at encoding temporally focused, task-relevant clinical signals. Notably, GTE-Qwen2-7B demonstrated greater robustness to extended inputs, highlighting its stronger capacity to extract salient information from longer contexts. Finally, combining LLM embeddings with CLMBR-T-Base representations improved overall performance, suggesting that these models capture complementary dimensions of patient history. This synergy may stem from CLMBR-T-Base’s limited code-based input window of 496 medical codes [13], or from the broader world knowledge encoded in LLMs that is absent in domain-specific models.
@@
Unlike count-based models and specialized EHR foundation models, LLM-based methods are agnostic to coding systems and data formats. Traditional EHR models rely on predefined vocabularies such as SNOMED, LOINC, and ICD, which can limit their applicability across healthcare settings. For instance, CLMBR-T-Base processes only 65,536 codes (26,249 unique), selected based on their entropy from the EHRSHOT dataset [31]. This restriction may lead to the exclusion of clinically important information when applied to external datasets with unseen codes. In contrast, LLMs operate on textual inputs, allowing them to interpret any clinical code mapped to a human-readable description. This flexibility is particularly valuable given the ongoing challenges in EHR interoperability due to varying coding practices, privacy constraints, and regulatory barriers, which often hinder the aggregation of large, standardized datasets for model pre-training [38]. Moreover, since LLMs are pre-trained on broad and diverse text corpora, including medical literature and case reports, they are well-equipped to capture the semantics of rare or underrepresented clinical concepts [34]. This mitigates the limitations of count-based models that often exclude infrequent events, ensuring that even rare but clinically significant phenomena are effectively encoded [39]. Finally, the ability to represent both structured and unstructured information makes LLMs well-suited for future multimodal EHR applications, including those that integrate clinical notes, diagnostic reports, and imaging metadata [40].
Our findings highlight that LLM embedding methods combine the strengths of count-based models, specialized EHR models, and the generalization capabilities of LLMs. While count-based and specialized EHR foundation models can be tuned to output well-calibrated probabilities, thereby enhancing trust in clinical applications, they are limited by predefined vocabularies and constrained training data. In contrast, LLMs can process any textual representation, enabling them to handle rare or unseen codes, though they often produce unstructured outputs that complicate clinical grounding and calibration [24]. LLM embedding methods bridge this gap by using the representational power of LLMs while maintaining compatibility with traditional predictive modeling frameworks. By converting EHR data into structured text and embedding it with general-purpose LLMs, we enable the use of standard classifiers such as logistic regression to generate clinically meaningful predictions. This approach retains the calibration strengths of traditional models while benefiting from the rich contextual understanding of LLMs.
This study has several limitations. First, our approach relies on a manually designed EHR serialization designed to capture medically relevant information, which may introduce bias when comparing against models trained on raw data. Additionally, LLM embedding model performance is sensitive to the specific content and instructions in the serialized text, which may limit reproducibility and generalization. While the models achieve strong predictive accuracy across tasks, including few-shot settings, their large parameter counts lead to longer computation time and higher resource demands. Moreover, the use of LLM embeddings requires training a separate downstream classifier, foregoing the native zero-shot or few-shot capabilities of LLMs. We also limited EHR serialization to 4,096 tokens due to runtime constraints, which may exclude relevant long-term historical data. Finally, evaluation was conducted on only two datasets, which may limit the generalizability of our findings across diverse healthcare systems.
@@
Future research should explore serialization-free approaches that allow LLMs to process raw EHR data directly, reducing biases introduced by manual text transformation. Integrating zero-shot and few-shot prompting into the LLM-based embedding framework could further enhance model flexibility and reduce dependency on downstream training. Extending the effective context window beyond 4,096 tokens will be important for capturing comprehensive patient histories. Additionally, developing techniques to distill large LLMs into smaller, more efficient models could improve their practical use in clinical settings. Expanding evaluations to real-world deployments and investigating how to combine the complementary strengths of domain-specific EHR models and general-purpose LLMs will be essential for building robust, scalable EHR foundation models.
@@ 4.1 EHRSHOT Database and Prediction Task
The EHR data used in our experiments is from the EHRSHOT benchmark for few-shot evaluation of EHR foundation models [31]. We obtained version 2.1 of the dataset, which is accessible via gated access under a research data use agreement. This dataset comprises longitudinal records for 6,739 patients, 921,499 visits, and 41,661,637 clinical events collected between 1990 and February 8, 2023. Each clinical event is linked to a specific patient and includes information such as start time, end time, a semantic code, a value, unit, visit ID, and the corresponding OMOP source table. We used the official ehrshot-benchmark repository111GitHub repository: https://github.com/som-shahlab/ehrshot-benchmark as a starting point to design our experiments, enabling us to build on existing functionalities and to facilitate comparisons with prior methods. The benchmark uses the Framework for Electronic Medical Records (FEMR) 222GitHub repository: https://github.com/som-shahlab/femr, which provides Python classes for efficient loading and processing of EHR data. All extensions and experiments conducted for this paper have been made publicly available via our GitHub repository: https://github.com/stefanhgm/ehrshot-benchmark. The EHRSHOT benchmark defines a rigorous evaluation including 15 clinical prediction tasks categorized into four groups: operational outcomes, anticipating lab test results, assignment of new diagnoses, and anticipating chest X-ray findings [31]. Task labels are derived from clinical events; hence, a single patient could contribute multiple labels per task, resulting in significant variations in task-specific sample sizes. For instance, frequent events like lab tests provide considerably more examples compared to rarer events such as new diagnoses. The benchmark focuses on analyzing model performance in a few-shot setting, which is particularly relevant for large pre-trained foundation models [14] due to their ability to generalize from limited training data. To this end, the benchmark defines evaluation settings with a constrained number of training and validation examples. Specifically, for k in 1, 2, 4, 8, 12, 16, 24, 32, 48, 64, 128, the benchmark uses k positive and k negative training examples, along with k positive and k negative validation examples, to train and tune supervised classifiers. Testing is always performed on the full set of examples. The classifiers evaluated within the EHRSHOT framework include logistic regression, random forests, and GBMs [41]. Performance is reported using the AUROC and the AUPRC. For few-shot settings, we average the results over five runs with different seeds and compute bootstrapped 95% confidence intervals [31]. Macro averages are reported for each task group, and an overall macro average is provided across all groups.

### v4

@@ 2.4 Effect of the EHR Serialization on LLM-Based Embedding Performance
| EHR Serialization Formats | |||||
| Markdown (ours) | |||||
| List events | |||||
| List events with time | |||||
| JSON | |||||
| XML | |||||
| YAML | |||||
To assess the impact of our hand-crafted EHR serialization, we compared it with a very simple newline-separated event list ordered from most recent to oldest, with and without timestamps. Overall performance decreased only slightly from 0.761 (0.733–0.788) to 0.751 (0.728–0.773) (with timestamps) and 0.751 (0.726–0.776) (without timestamps). Task-group performance shifted, with gains for operational outcomes, assignment of new diagnoses, and chest X-ray findings, likely reflecting different information prioritization in the serialization. Thus, while a hand-crafted format is beneficial, the embedding method is not highly sensitive to very simple data formats. Two variants with reversed ordering performed substantially worse (see Table 11). Replacing Markdown with JSON, XML, or YAML yielded minimal differences (see Table 5). XML showed the largest decrease, which might be due to its greater formatting overhead reducing effective clinical content.
We further quantified the contribution of individual serialization components with ablations using the Qwen3-Emb-8B model (see Fig. 6 and Table 12). For these runs we used a 4,096-token limit due to computational constraints. Replacing task-specific instructions with a generic prompt slightly reduced performance, and removing instructions led to a further decrease. The largest performance drop was observed for lab-test prediction, indicating that task-aligned instructions help the model attend to relevant spans.
Removing individual components from the full serialization generally had modest effects, suggesting redundancy across data sources. A notable exception was aggregated lab results, whose removal caused a large drop in lab-test prediction from 0.866 (0.859–0.873) to 0.719 (0.706–0.731), underscoring dependence on recent lab results. Interestingly, assignment-of-new-diagnoses performance improved when condition histories were omitted, driven primarily by celiac disease (0.630 to 0.734) and lupus (0.680 to 0.743), suggesting that long lists of past diagnoses can add noise for difficult-to-ascertain or miscoded conditions.
In serializations consisting of a single component, no component matched the full EHR, underscoring the value of combining modalities. The strongest single component was conditions, achieving an AUROC of 0.707 (0.679–0.736), followed by lab results, medications, and procedures. For lab-test prediction specifically, a lab-only serialization outperformed the full EHR, likely because unrelated context was reduced. Demographics, body metrics, vital signs, and visits yielded much lower overall macro performance (see Table 12). Yet, each showed task-specific utility: demographics aided new-diagnosis assignment, body metrics supported operational outcomes, and vital signs and visits were more informative for chest X-ray findings (see Fig. 6).
2.5 Effect of Context Length and Temporal Scope on Embedding Effectiveness
To assess model sensitivity to input length and recency, we performed ablation studies on temporal window and context size using EHRSHOT data (see Fig. 7 and Table 13). The models differed in optimal context sizes, ranging from 8,192 tokens for Qwen3-Emb-8B, 4,096 for Qwen2-Emb-7B, and 2,048 for Llama3.1-LLM2Vec-8B. This pattern suggests that newer LLM embedding architectures are better able to handle longer contexts. All models exhibited sharp performance drops at 512 tokens, a common constraint in older language models [49]. All LLM embedding models improved when restricted to more recent data, with the 1-month window outperforming longer histories. Qwen3-Emb-8B and Qwen2-Emb-7B even slightly outperformed CLMBR-T-Base with a macro-AUROC of 0.773 (0.746–0.800) and 0.773 (0.746–0.801) (see Table 14) compared to 0.769 (0.746–0.792) for CLMBR-T-Base using the full data (see Table 3). Because CLMBR-T-Base is restricted to 496 input codes, it effectively encodes a limited time horizon. To disentangle long-context capacity from effects of time windows, we additionally evaluated the count-based baseline across the same temporal windows. It followed the same recency trend, indicating an inherent difficulty in performing prediction with historic data in the EHRSHOT benchmark. Using data from the day before prediction remained strong across methods, whereas restricting to the last hour led to a pronounced decline.
2.6 Fine-Tuned LLM Embedding and LLM Decoder Models
To further contextualize the use of general-purpose LLMs for clinical prediction, we evaluate decoder-style LLMs that generate text, where token probabilities can be used for classification. We extend the prompt with an additional instruction to produce Yes and No tokens leveraged for prediction (see Table 9). Because of the length of the serialized EHRs, in-context few-shot prompting was infeasible, and the decoder was evaluated only in the zero-shot prompting setting. In addition, we fine-tuned both the LLM embedding model and the decoder using Low-Rank Adaptation (LoRA). The decoder experiments train the instruction-tuned causal language model Qwen3-8B to generate direct Yes and No predictions for each serialized EHR, while the encoder experiments pair Qwen3-Emb-8B with a supervised classification head (see Section 4.6). Both approaches preserve the EHRSHOT task definitions, operate on identical Markdown-based EHR serializations, and employ the same few-shot sampling strategy across all shot sizes and replicates. Due to computational constraints, we used a token limit of 4,096 tokens, and for the Anticipating Lab Test Results task group we only used a uniformly random subset of 4,096 test samples.
| Model | Operational Outcomes | Anticipating Lab Test Results | Assignment of New Diagnoses | Anticipating Chest X-ray Findings | Macro Avg. Across Task Groups |
| Encoder models | |||||
| Qwen3-Emb-8B | |||||
| Qwen3-Emb-8B (LoRA) | |||||
| Decoder models | |||||
| Qwen3-8B (Zero-Shot) | |||||
| Qwen3-8B (LoRA) | |||||
The zero-shot Qwen3-8B decoder achieves a macro-AUROC of 0.651 (0.593–0.707), outperforming all other model variants for up to eight training examples per class (see Fig. 8). This pattern reflects architectural differences between the two approaches: the decoder can directly exploit pretrained logits for Yes and No tokens, whereas embedding variants must learn a classification head from minimal supervision. At , the fine-tuned decoder surpasses its zero-shot counterpart (0.659 (0.603–0.714)) and continues to improve with additional data (0.707 (0.653–0.758) at , 0.738 (0.684–0.788) at ).
The encoder-based LLM embedding model exhibits a complementary trade-off. Fine-tuning Qwen3-Emb-8B yields steady improvements (0.526 (0.461–0.590) at to 0.716 (0.657–0.770) at ), yet the LoRA-tuned encoder consistently underperforms the frozen-encoder baseline reported in the main results across all shot sizes. We expect that larger shot counts or refined hyperparameters could narrow this gap, though computational limits precluded exploration of higher or extensive hyperparameter sweeps. At , the fully fine-tuned Qwen3-8B decoder achieves a macro-AUROC of 0.738 (0.684–0.788), marginally exceeding the fine-tuned embedding model (0.727 (0.676–0.775)), but this modest gain comes at substantially higher optimization and inference costs.
@@ 3 Discussion
Our study shows that general-purpose LLM embedding models, pretrained on large-scale natural language corpora, can be effectively repurposed for clinical prediction using EHR data. The model Qwen3-Emb-8B outperformed a strong count-based baseline and, in several clinical domains, matched or surpassed the dedicated EHR foundation model CLMBR-T-Base [13, 44]. This is notable given that CLMBR-T-Base was trained on data from the same health system as EHRSHOT. Beyond in-domain performance, the strong results on the UKB dataset highlight the generalization capabilities of LLM-based encoders under population and coding shifts. Across a broad set of tasks, including few-shot settings, our findings indicate that knowledge acquired during large-scale text pretraining enables LLMs to extract clinically meaningful patterns from serialized EHRs. Performance scaled with model size and newer architectures, suggesting further gains are likely with future LLM development. In contrast, domain-specific EHR models remain constrained by the scarcity and fragmentation of large, diverse clinical datasets. Overall, these results support general-purpose LLMs as scalable, adaptable, and high-performing alternatives for EHR representation and clinical prediction.
Our experiments suggest several key factors contribute to the success of LLM embedding models for encoding EHR data. First, task-specific instructions prepended to the EHR serialization improved performance, especially for targets that require focusing on particular spans such as lab results, aligning with the instruction-tuned nature of modern embedding models [39, 42]. Second, aggregating time series data was critical: removing aggregated laboratory values produced a large drop for lab-test prediction, underscoring the relevance of temporally proximate quantitative measurements. At the same time, the approach was robust to the choice of serialization: a very simple newline-separated event list performed only slightly worse than our hand-crafted Markdown serialization. Also, alternative data formats such as JSON and YAML showed comparable results, suggesting that model capacity rather than markup specifics drives most of the prediction performance. Third, temporal scope and context length mattered. Intermediate time windows, for example one month prior to prediction, were consistently optimal across models, indicating that recent history carries disproportionate predictive value. Long input contexts of up to 8,192 tokens could only be used effectively by Qwen3-Emb-8B, while older LLM embedding models showed a decrease in performance. Finally, we observed practical trade-offs between single-embedding serializations and multi-embedding variants like MEME [34]. While MEME produced modest average gains for Qwen3-Emb-8B on EHRSHOT, it required roughly eight times more model inferences, raising cost and latency considerations in practice.
@@
Unlike count-based models and specialized EHR foundation models, LLM-based methods are agnostic to coding systems and data formats. Traditional EHR models rely on predefined vocabularies such as SNOMED, LOINC, and ICD, which can limit applicability across healthcare settings. For instance, CLMBR-T-Base processes 26,249 unique codes, selected from EHRSHOT [44]. When applying CLMBR-T-Base to UKB, only 7,969 of 50,702 unique codes, 16%, mapped to this fixed vocabulary, rendering most codes invisible at inference. In contrast, LLM encoders operate on text, allowing any clinical code to be incorporated via a human-readable description at serialization time. This flexibility directly addresses interoperability challenges arising from site-specific coding practices and privacy constraints that hinder pooled pretraining [50]. Moreover, broad text pretraining, often including biomedical sources, helps capture the semantics of rare or underrepresented concepts [45], mitigating the bias of count-based approaches that downweight infrequent events [51]. The same textual interface naturally supports future multimodal extensions that integrate notes, diagnostic reports, and imaging metadata [52].
Our findings highlight that LLM embedding methods combine the strengths of count-based models, specialized EHR models, and the generalization capabilities of LLMs. Count-based and specialized EHR foundation models can yield well-calibrated probabilities but are constrained by fixed vocabularies and local training distributions. Pure generation with LLMs can struggle with grounding and calibration [26]. Embedding-based pipelines bridge this gap: they leverage the representational power of LLMs while remaining compatible with standard discriminative heads such as logistic regression. This approach retains familiar calibration workflows and enables straightforward ensembling with domain-specific representations, as shown by the Qwen3-Emb-8B + CLMBR-T-Base concatenation.
We compared frozen LLM embeddings, a decoder-style LLM that directly emits Yes and No tokens, and parameter-efficiently fine-tuned variants of both. Two key observations emerged. First, the zero-shot decoder approach was competitive in the very low-shot regime for up to eight training examples even without task-specific training, likely because pretrained logits for Yes and No tokens can be exploited immediately. With more supervision for , LoRA-tuned decoding further improved, reaching a macro-AUROC of 0.738 at . Second, end-to-end LoRA fine-tuning of Qwen3-Emb-8B with a linear head produced steady gains with increasing , but did not surpass the frozen-encoder baseline within the explored budgets. This supports using a frozen embedding model with a tunable classification head for LLM embedding–based EHR prediction [44]. Leveraging token outputs from decoder-style LLMs can be a viable alternative when a single, well-defined task justifies the higher training and inference costs, or when strong very-few-shot performance is required but in-context examples are infeasible due to long inputs.
This study has several limitations. First, our approach relies on a manually designed EHR serialization designed to capture medically relevant information; while ablations suggest the method is robust to simpler formats, hand-crafted choices can still introduce bias relative to models trained on raw, structured data. Second, LLM embedding model performance is sensitive to the specific content and instructions in the serialized text, which may limit reproducibility and generalization. Third, LLM encoders have larger computational footprints than CLMBR-T-Base for population-scale screening, and many analyses were constrained to 4,096–8,192 tokens for runtime reasons, potentially excluding long-term history. Fourth, we evaluated two large cohorts, EHRSHOT and UKB; broader external validation, including settings with different documentation practices, laboratory availability, and care pathways, will be necessary to assess generalizability, calibration, and fairness at scale. Finally, while our decoder experiments provide a complementary perspective, we did not exhaust the fine-tuning design space, for example data scaling, multi-task objectives, pooling strategies, or calibration tuning, and in-context learning was impractical because serialized EHRs typically filled the context window.
@@
Future research should explore serialization-free approaches that allow LLMs to process raw EHR tables and timelines, reducing bias from manual text transformation. Improved pooling or attention-on-attention mechanisms over long inputs may preserve salient signals without diluting them at extreme context lengths. Extending the effective context window beyond 8,192 tokens will be important for capturing comprehensive patient histories. Additionally, developing techniques to distill large LLMs into smaller, more efficient models could improve their practical use in clinical settings. Finally, expanding evaluations to real-world deployments and investigating how to combine the complementary strengths of domain-specific EHR models and general-purpose LLMs will be essential for building robust, scalable EHR foundation models.
@@ 4.1 EHRSHOT Database and Prediction Task
The EHR data used in our experiments is from the EHRSHOT benchmark for few-shot evaluation of EHR foundation models [44]. We obtained version 2.1 of the dataset, which is accessible via gated access under a research data use agreement. This dataset comprises longitudinal records for 6,739 patients, 921,499 visits, and 41,661,637 clinical events collected between 1990 and February 8, 2023. Each clinical event is linked to a specific patient and includes information such as start time, end time, a semantic code, a value, unit, visit ID, and the corresponding OMOP source table. We used the official ehrshot-benchmark repository111GitHub repository: https://github.com/som-shahlab/ehrshot-benchmark as a starting point to design our experiments, enabling us to build on existing functionalities and to facilitate comparisons with prior methods. The benchmark uses the Framework for Electronic Medical Records (FEMR) 222GitHub repository: https://github.com/som-shahlab/femr, which provides Python classes for efficient loading and processing of EHR data. All extensions and experiments conducted for this paper are publicly available via our GitHub repository: https://github.com/stefanhgm/ehrshot-benchmark. The EHRSHOT benchmark defines a rigorous evaluation including 15 clinical prediction tasks categorized into four groups: operational outcomes, anticipating lab-test results, assignment of new diagnoses, and anticipating chest X-ray findings [44]. Task labels are derived from clinical events, so a single patient can contribute multiple labels per task, resulting in substantial variation in task-specific sample sizes. For instance, frequent events like laboratory tests yield many more examples than rarer events such as incident diagnoses. The benchmark focuses on analyzing model performance in a few-shot setting, which is particularly relevant for large pretrained foundation models [14]. Specifically, for , the protocol uses positive and negative training examples and positive and negative validation examples to train and tune supervised classifiers. Testing is always performed on the full set of available examples for each task. The classifiers evaluated within the EHRSHOT framework include logistic regression, random forests, and GBMs [53]. Performance is reported using the AUROC and the AUPRC. For few-shot settings, we average the results over five runs with different seeds and compute bootstrapped 95% confidence intervals [44]. Mean averages are reported for each task group, and an overall macro average is provided across all groups.